# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt (eccv) Package eccv Warning: Package ‘hyperref’ is loaded with option ‘pagebackref’, which is *not* recommended for camera-ready version

### v1

@@ 5 Conclusion
In this paper, we propose a mask classification prompt-free SAM adaptation framework for medical image segmentation, named MaskSAM, which adapts pre-trained SAM models worked on from 2D natural images to 3D medical images without any prompts provided. By designing a prompt generator combined with the image encoder in SAM to generate a set of auxiliary classifier tokens, auxiliary binary masks, and auxiliary bounding boxes. Each pair of auxiliary mask and box prompts, which can solve the requirements of extra prompts, is associated with class label predictions by the sum of the auxiliary classifier token and the learnable global classifier tokens in the mask decoder of SAM to solve the predictions of semantic labels. By inserting one of our redesigned 3D depth-convolution adapter (DConvAdapter) for image embeddings and 3D depth-MLP adapter (DMLPAdapter) for prompt embeddings into each transformer block in the image encoder and mask decoder, our model enables pre-trained 2D SAM models to extract 3D information and adapt to 3D medical images. Our method achieves state-of-the-art performance on AMOS2022 [27], 90.52% Dice, which improved by 2.7% compared to nnUNet. Meanwhile, our method surpasses nnUNet by 1.7% on ACDC [3] and 1.0% on Synapse [29] datasets.
Appendix 0.A Implementation Details
Our models are based on the codebase of nnUNet for the preprocessing and postprocessing. For the preprocessing, we utilize some data augmentations such as rotation, scaling, Gaussian noise, Gaussian blur, brightness, and contrast adjustment, simulation of low resolution, gamma augmentation, and mirroring. During training, we set the initial learning rate to 0.01 and employ a “poly” decay strategy in Eq. (1).
$lr(e)=init\_lr\times(1-\frac{e}{\rm MAX\_EPOCH})^{0.9},$ (1)
where $e$ means the number of epochs, MAX_EPOCH means the maximum of epochs, set it to 500, 1000, and 1000 for ACDC, AMOS2022, and Synapse dataset respectively, and each epoch includes 250 iterations. We utilize SGD as our optimizer and set the momentum to 0.99. The weighted decay is set to 3e-5. For the postprocessing, we crop the whole image into several patches with half overlapping. For each patch, our model infers 8 times by different three axes (i.e. axial, sagittal, and coronal planes) and then averages all output to get the final predictions. All experiments are conducted using two NVIDIA RTX A6000 GPUs with 40GB memory with a batch size of 2.
Method Spleen R.Kd L.Kd GB Eso. Liver Stomach Aorta IVC Veins Pancreas AG Average
TransUNet 95.2 92.7 92.9 66.2 75.7 96.9 88.9 92.0 83.3 79.1 77.5 63.7 83.8
nnUNet 94.2 89.4 91.0 70.4 72.3 94.8 82.4 87.7 78.2 72.0 68.0 61.6 80.2
3D UX-Net 94.6 94.2 94.3 59.3 72.2 96.4 73.4 87.2 84.9 72.2 80.9 67.1 81.4
SwinUNETR 95.6 94.2 94.3 63.6 75.5 96.6 79.2 89.9 83.7 75.0 82.2 67.3 83.1
nnFormer 93.5 94.9 95.0 64.1 79.5 96.8 90.1 89.7 85.9 77.8 85.6 73.9 85.6
SAMed_h 95.3 92.1 92.9 62.1 75.3 96.4 90.2 87.6 79.8 74.2 77.9 61.0 82.1
MaskSAM (Ours) 97.0 95.3 95.3 68.6 78.3 97.3 92.4 91.8 87.6 80.4 86.9 72.3 86.9
Table 1: Quantitative results on BTCV dataset.
Appendix 0.B Theoretical comparison with SAM-based models.
The main contributions of our MaskSAM different from the existing SAM-based models are i) automatic prompt, ii) the classifier to generate semantic labels for each mask, and iii) remain all parameters of the original SAM for zero-shot capabilities. There are several categories of the existing SAM-based models. The first category does not modify SAM, such as MedSAM and Polyp-SAM. These models need manual prompts, such as points or boxes, and cannot classify masks into semantic labels. The second category uses parameter-efficient transfer learning, such as Adapters, into SAM. The popular model, Med-SA, uses the GT to generate prompts during inference, which do not have any practical clinical values. It also includes the non-automatic models of the 3DSAM-Adapter and MA-SAM. These models do not handle the requirements of extra prompts. The third category is that cannot classify masks into semantic labels, such as DeSAM, Med-SA, and MA-SAM. Since SAM only predicts binary masks, these models do not address the lack of classifiers. The fourth category is abandoning the components of SAM, such as Mask Decoder, to handle the inability to classify semantic labels, such as 3DSAM-Adapter. This way inevitably destroys the consistency and zero-shot capabilities of SAM. These models only use the pre-trained ViT encoder, which is not the contribution of SAM.
Appendix 0.C More experiments.
We conducted the experiments on the Beyond the Cranial Vault (BTCV) challenge datase again but different split strategies and the Pancreas Tumor Segmentation task within 2018 MICCAI Medical Segmentation Decathlon Challenge (MSD-Pancreas) dataset. We follow the split in SwinUNETR which contains 24 cases for training and 6 cases for testing for BTCV and the results are shown in Tab. 2. We follow the split in 3DSAM-Adapter that the datasets are randomly split into 70%, 10%, and 20% for training, validation, and testing and the results are shown in Tab. 1. Our model achieves the state-of-the-art performance and outperforms nnUNet by 6.7% and 1.0% on BTCV and MSD-Pancreas dataset. Meanwhile, our model surpasses the SAM-based methods, SAMed_h by 4.9% on BTCV, MA-SAM by 2.4% on MSD-Pancreas and 3DSAM-Adapter by 12.4% on MSD-Pancreas dataset. In conclusion, the effectiveness of our method is robustly demonstrated.
Methods DSC $\uparrow$ NSD $\uparrow$
nnUNet 41.6 62.5
3D UX-Net 34.8 52.6
SwinUNETR 40.6 60.0
nnFormer 36.5 54.0
3DSAM-Adapter 30.2 45.4
MA-SAM 40.2 59.1
MaskSAM (Ours) 42.6 68.6
Table 2: Quantitative results on MSD-Pancreas dataset.

### v2

@@ 4 Conclusion
In this paper, we propose a mask classification prompt-free SAM adaptation framework for medical image segmentation, named MaskSAM. By designing a prompt generator combined with the image encoder in SAM to generate a set of auxiliary classifier tokens, binary masks, and bounding boxes to solve the inability to infer semantic labels and its reliance on providing extra prompts. By inserting one of our redesigned 3D depth-convolution adapter (DConvAdapter) for image embeddings and 3D depth-MLP adapter (DMLPAdapter) for prompt embeddings into each transformer block in SAM for efficient fine-tuning, our model enables pre-trained 2D SAM models to extract 3D information and adapt to 3D medical images. Our method achieves state-of-the-art performance on AMOS2022 [25], ACDC [3], and Synapse [27] datasets with 90.52%, 93.39%, and 87.23% Dice, which improved by 2.7%, 1.7%, and 1.0% compared to nnUNet, respectively.